Jordan Jones (JJ) and Casey Carter (CC) are portfolio managers at your firm. Each manages a well-diversified portfolio. Your boss has asked for your opinion regarding their performance in the past year. JJ’s portfolio has a beta of 0.6 and had a return of 8.5%; CC’s portfolio has a beta of 1.4 and had a return of 9.5%. Which manager had better performance? Why?
Answer #1

Let the Risk Free Rate (Rf) is 5%

Treners Ratio = (Return of Portfolio - Rf) / Beta

Treners Ratio of JJ = (0.085- 0.05) / 0.6 = 0.05833

Treners Ratio of CC = (0.095- 0.05) / 1.4 = 0.03214

Treners Ratio of JJ is better.

Return per risk (beta) for JJ is better than CC.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.